OpenStudy (anonymous):

Solve for a A=(1/2)h(a+b) I have 2A=ah+bh help please?

OpenStudy (anonymous):

2A = ah + bh then subtract bh from both sides -bh -bh 2A - bh = ah then divide both sides by h /h /h /h so you get a = (2A)/h - b

OpenStudy (anonymous):

Thank you! That helped.